Suppose that the market demand and the market supply for a smart phone are respectively given by P space equals space 500 space minus space 0.01 Q to the power of d and P space equals space 200 plus thin space 0.005 space Q to the power of s . What is the equilibrium price?
1
Expert's answer
2020-10-05T13:47:39-0400

p=500-0.01Qd

Qd=(500-p)/0.01

Qd=50,000-100p


p=200+0.005Qs

Qs=(p-200)/0.005

Qs=200p-40,000


Qd=Qs

50,000-100p=200p-40,000


p=300